OpenStudy (sarahc):

A certain species of tree grows an average of 4.2 cm per week. Write an equation for the sequence that represents the weekly height of this tree in centimeters if the measurements begin when the tree is 300 centimeters tall.

OpenStudy (mathstudent55):

At week zero, the height of the tree is 300 cm. That means that for x = 0, y = 300 Each week, x, the tree grows 4.2 cm. The total height at any week is the beginning height (300 cm) plus the weekly growth (4.2x). y = 4.2x + 300 Just to check, let's look at two points. At week 0, the height should be 300 cm. For x = 0, y = 4.2 * 0 + 300 y = 300 (This checks out) After 1 week, the height should be 304.2 cm. That means we need to let x = 1 (1 week of growth) and check the height. For x = 1, y = 4.2 * 1 + 300 y = 4.2 + 300 y = 304.2 (This also checks out) Our equation is correct: y = 4.2x + 300

OpenStudy (sarahc):

@mathstudent55 thank you!! do u mind checking out one more?

OpenStudy (mathstudent55):

You're welcome.